OpenStudy (anonymous):

Please factor the following: x^3y^2+xy^5

OpenStudy (anonymous):

i would pull out the y^2 from it you basically divide both by y^2 and make y^2 the coefficient for the whole term: x3y2+xy5= y2(x3+xy3) then pull out an x and do the same as you did with y y2(x3+xy3)= xy^2(x^2+y^3)

OpenStudy (anonymous):

Thanks for the help!